please help me, would really appreciate your help! =)
P^3+2p^2-8p=0

Answer by ewatrrr(24785)


Hi
solving:
p^3+2p^2-8p=0
p(p^2+2p-8)=0
p = 0
(p^2+2p-8)=0
factorin
(p+4)(p-2)= 0 Note:SUM of the inner product(4p) and the outer product(-8p) = 2p
(p+4)= 0 p = -4
(p-2)= 0 p = 2